OpenStudy (anonymous):

Solve the equation. (Enter all answers including repetitions. If there is no solution, enter NO SOLUTION.) sqrt 20-x + sqrt20+x=8 x= (smaller value) x= (larger value)

myininaya (myininaya):

i suppose you mean \[\sqrt{20-x}+\sqrt{20+x}=8\]

myininaya (myininaya):

is that right? its not really what you have written but sometimes i have esp

OpenStudy (anonymous):

yes!

myininaya (myininaya):

ok subtract want those square root thingys on both sides and then square and then you will have to get another square rooty by itself and then square to get rid of that one

myininaya (myininaya):

\[\sqrt{20-x}=8-\sqrt{20+x}\] \[(\sqrt{20-x)}^2=(8-\sqrt{20+x})^2\] \[20-x=8^2-2*8*\sqrt{20+x}+(20+x)\]

myininaya (myininaya):

let me know if if you still are having trouble with this

OpenStudy (anonymous):

I just multiply?

myininaya (myininaya):

\[20-x=64-16\sqrt{20+x}+20+x\]

myininaya (myininaya):

move everything over except the square root thingy

myininaya (myininaya):

if you want to leave the 16 with it you can

myininaya (myininaya):

\[20-20-x+x-64=-16\sqrt{20+x}\] \[-64=-16\sqrt{20+x}\] \[64=16\sqrt{20+x}\] \[\frac{64}{16}=\frac{16\sqrt{20+x}}{16}\] \[4=\sqrt{20+x}\]

OpenStudy (anonymous):

;S

myininaya (myininaya):

whats that mean

OpenStudy (zarkon):

i see a small error

myininaya (myininaya):

uh oh

OpenStudy (zarkon):

with your x's

OpenStudy (zarkon):

they don't cancel

myininaya (myininaya):

i see it too

myininaya (myininaya):

i subtracted it on one sides and added it on the other lol

myininaya (myininaya):

\[-64-2x=-16\sqrt{20+x}\]

OpenStudy (zarkon):

You shouldn't be doing this stuff this late at night...not good for the brain ;)

myininaya (myininaya):

lol

myininaya (myininaya):

\[-(64+2x)=-16\sqrt{20+x}\] \[64+2x=16\sqrt{20+x}\]

myininaya (myininaya):

\[(64+2x)^2=(16\sqrt{20+x})^2\] \[64^2+2*64*2x+(2x)^2=16(20+x)\] \[4096+256x+4x^2=320+16x\] \[4x^2+256x-16x+4096-320=0\] \[4x^2+240x+3776=0\] please let me know if made anymore mistakes

myininaya (myininaya):

you can divide both sides by 4

myininaya (myininaya):

then you might want to try factoring or you can use the quad formula

myininaya (myininaya):

check your solutions you may get one that does not work

myininaya (myininaya):

or both

OpenStudy (zarkon):

you missed the \[16^2\] on the rhs

myininaya (myininaya):

no!!!!!!!!!!

myininaya (myininaya):

i cant do it anymore

OpenStudy (zarkon):

lol

myininaya (myininaya):

deep breath! \[64^2+2*64*2x+(2x)^2=16^2(20+x)\]

OpenStudy (zarkon):

lol

myininaya (myininaya):

\[4096+256x+4x^2=256(20+x)\] \[4096+256x+4x^2=5120+256x\] \[4x^2+256x-256x=5120-4096\]

myininaya (myininaya):

\[4x^2=1024\]

myininaya (myininaya):

this is easy to solve then the wrong equation i gave you earlier lol

OpenStudy (zarkon):

nice

myininaya (myininaya):

please tell me you can finish this ep4710 thanks for your help zarkon

OpenStudy (zarkon):

you deserve another medal for all that work

myininaya (myininaya):

omg i suck

myininaya (myininaya):

too many errors

OpenStudy (zarkon):

na

OpenStudy (zarkon):

it's all good

myininaya (myininaya):

you know your students are learning if they are making corrections to your work i just hope people notice them not you zarkon because you know everything but yeah he probably wouldn't have caught them